# Study of semileptonic $B\to DP\ell^{+}\nu_{\ell}$ decays based on the SU(3) flavor symmetry

### v1

@@ III Decays $B\to DP\ell^{+}\nu_{\ell}$ with the $D^{*},D_{0},D^{*}_{2}$ resonances
For the $B\to DP\ell^{+}\nu_{\ell}$ decays with the resonances, in the case of the decay widths of the resonance states are very narrow, the resonant branching ratios respect a simple factorization relation
@@
where $D_{J}=D^{*},D_{0},D^{*}_{2}$. And this result is also a good approximation for wider resonances. Eq. (28) will be used in our analysis for resonant $B\rightarrow D_{J}(\rightarrow DP)\ell^{+}\nu_{\ell}$ decays. Relevant $\mathcal{B}(B\rightarrow D_{J}\ell^{+}\nu_{\ell})$ and $\mathcal{B}(D_{J}\rightarrow DP)$ also can be obtained by the SU(3) flavor symmetry.
III.1 Semileptonic three-body $B\to D_{J}\ell^{+}\nu_{\ell}$ decays
The helicity amplitudes of the decays $B\to D_{J}\ell^{+}\nu_{\ell}$ can be written as
$\mathcal{M}(B\to D_{J}\ell^{+}\nu_{\ell})$ $=$ $\frac{G_{F}}{\sqrt{2}}V_{cb}\sum_{mm^{\prime}}g_{mm^{\prime}}L^{\lambda_{\ell}\lambda_{\nu}}_{m}H^{\lambda_{M}}_{m^{\prime}},$ (29)
@@ III.1 Semileptonic three-body $B\to D_{J}\ell^{+}\nu_{\ell}$ decays
$L^{\lambda_{\ell}\lambda_{\nu}}_{m}$ $=$ $\epsilon_{\alpha}(m)\bar{\nu_{\ell}}\gamma^{\alpha}(1-\gamma_{5})\ell,$ (30)
$H^{\lambda_{D_{J}}}_{m^{\prime}}$ $=$ $\epsilon^{*}_{\beta}(m^{\prime})\langle{D_{J}}(p,\epsilon^{*})|\bar{c}\gamma^{\beta}(1-\gamma_{5})b|B(p_{B})\rangle,$ (31)
where the particle helicities $\lambda_{D_{J}}=0$ for $D/D^{*}_{0}$, $\lambda_{D_{J}}=0,\pm 1$ for ${D_{J}}=D^{*}/D^{*}_{2}$, $\lambda_{\ell}=\pm\frac{1}{2}$, $\lambda_{\nu}=+\frac{1}{2}$, and $\epsilon_{\mu}(m)$ is the polarization vectors of the virtual $W$ with $m=0,t,\pm 1$.
The matrix elements can be parameterized via the form factors of the $B\to D_{J}$ transitions Cheng:2017pcq ; Cheng:2003sm
$D^{*}_{0}(p)\left|\bar{c}\gamma_{\mu}\gamma_{5}b\right|B(p_{B})\right>$ $=$ $-i\Bigg{[}f_{1}^{BD_{0}}(q^{2})\Big{(}(p+p_{B})_{\mu}-\frac{m_{B}^{2}-m_{D_{0}}^{2}}{q^{2}}q_{\mu}\Big{)}+f_{0}^{BD_{0}}(q^{2})\frac{m_{B}^{2}-m_{D_{0}}^{2}}{q^{2}}q_{\mu}\Bigg{]},$ (33)
$+iq_{\mu}(\varepsilon^{*}.p_{B})\frac{2m_{D^{*}}}{q^{2}}[A^{BD^{*}}_{3}(q^{2})-A^{BD^{*}}_{0}(q^{2})],$
$D_{2}(p,\varepsilon^{*})\left|\bar{c}\gamma_{\mu}(1-\gamma_{5})b\right|B(p_{B})\right>$ $=$ $\frac{2iV^{BD_{2}}(q^{2})}{m_{B}+m_{D_{2}}}\epsilon_{\mu\nu\alpha\beta}e^{*\nu}p^{\alpha}_{B}p^{\beta}$ (35)
$+2m_{D_{2}}\frac{e^{*}\cdot q}{q^{2}}q_{\mu}A^{BD_{2}}_{0}(q^{2})+(m_{B}+m_{D_{2}})\Big{(}e^{*}_{\mu}-\frac{e^{*}\cdot q}{q^{2}}q_{\mu}\Big{)}A^{BD_{2}}_{1}(q^{2})$
$-\frac{e^{*}\cdot q}{m_{B}+m_{D_{2}}}\Big{(}(p_{B}+p)_{\mu}-\frac{m_{B}^{2}-m^{2}_{D_{2}}}{q^{2}}q_{\mu}\Big{)}A^{BD_{2}}_{2}(q^{2}),$
where $s=q^{2}$ ($q=p_{B}-p_{M}$), $\varepsilon^{*}$ is the polarization of vector meson, and $e^{*\nu}\equiv\frac{\varepsilon^{*\mu\nu}\cdot p_{B\mu}}{m_{B}}$, $f^{BD}_{+,0}$, $f^{BD_{0}}_{1,0}$
The hadronic helicity amplitudes can be written as
@@
$H^{D^{*}_{2}}_{\pm}$ $=$ $\frac{2|\vec{p}_{D^{*}_{2}}|}{2\sqrt{2}m_{D^{*}_{2}}}\left[(m_{B}+m_{D^{*}_{2}})A^{BD^{*}_{2}}_{1}(q^{2})\mp\frac{2m_{B}|\vec{p}_{D^{*}_{2}}|}{(m_{B}+m_{D^{*}_{2}})}V^{BD^{*}_{2}}(q^{2})\right],$ (41)
@@
for $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ decays, where $|\vec{p}_{D_{J}}|\equiv\sqrt{\lambda(m_{B}^{2},m_{D_{J}}^{2},q^{2})}/2m_{B}$ with $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2ab-2ac-2bc$.
The differential branching ratios of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays are Ivanov:2019nqd
with
${\cal H}_{U}$ $=$ $|H^{D_{J}}_{+}|^{2}+|H^{D_{J}}_{-}|^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ {\cal H}_{L}=|H^{D_{J}}_{0}|^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ {\cal H}_{S}=|H^{D_{J}}_{t}|^{2},$ (46)
where $\lambda\equiv\lambda(m_{B}^{2},m_{D_{J}}^{2},q^{2})$ and $m_{\ell}^{2}\leq q^{2}\leq(m_{B}-m_{D_{J}})^{2}$.
Now one can obtain the branching ratios of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays by the relevant form factors, which depend on the different methods. In this work, we use the SU(3) flavor symmetry to obtain the relations of the hadronic helicity amplitudes, and the same relations are also true for the form factors. In terms of the SU(3) flavor symmetry, the hadronic helicity amplitudes can be parameterized as
$H^{\lambda_{D_{J}}}_{m^{\prime}}=C^{D_{J}}_{0}B_{i}(D_{J})^{i},$ (47)
where $C^{D_{J}}_{0}$ are the nonperturbative coefficients of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays under the SU(3) flavor symmetry. For the charmed four-quark states $D_{0}^{4q}$, $H^{\lambda_{D_{0}}}_{m^{\prime}}=C^{4q,D_{0}}_{0}B_{i}(D_{J})^{ij}_{j}$. And the hadronic helicity amplitude relations for the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays are summarized in Tab. 5.
Table 3: SU(3) IRA amplitudes for $B\to D_{J}\ell^{+}\nu$ decays due to $\bar{b}\to\bar{c}\ell^{+}\nu$.
@@
Table 4: The experimental data and the SU(3) flavor symmetry predictions of the branching ratios of $B\to D/D^{*}\ell^{+}\nu_{\ell}$ decays within $2\sigma$ errors. $\mathcal{B}(B\to D/D^{*}\ell^{+}\nu_{\ell})$ is in unit of $10^{-2}$. adenotes that the experimental data are not used to constrain the parameter $C^{D^{*}}_{0}$.
@@
Table 5: The experimental data and the SU(3) flavor symmetry predictions of the branching ratios of the $B\to D_{0}/D_{2}^{*}\ell^{+}\nu_{\ell}$ decays within $2\sigma$ errors. 2q(4q)denote the two(four) quark state predictions. $\mathcal{B}(B\to D_{0}/D^{*}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ is in unit of $10^{-3}$, and $\mathcal{B}(B\to D_{0}/D^{*}_{2}\tau^{+}\nu_{\tau})$ is in unit of $10^{-4}$.
Branching ratios Exp. data Belle-II:2022evt Our predictions Branching ratios Exp. data Our predictions
$\mathcal{B}(B^{+}\to\overline{D}^{0}_{0}\ell^{\prime+}\nu_{\ell^{\prime}})$ $4.2\pm 1.6$ $3.98\pm 1.30$ $\mathcal{B}(B^{+}\to\overline{D}^{0}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ $2.9\pm 0.6$ Belle-II:2022evt $3.20\pm 0.30$
$\mathcal{B}(B^{0}\to D^{-}_{0}\ell^{\prime+}\nu_{\ell^{\prime}})$ $3.9\pm 1.4$ $3.71\pm 1.21$ $\mathcal{B}(B^{0}\to D^{-}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ $2.7\pm 0.6$ Belle-II:2022evt $2.99\pm 0.29$
$\mathcal{B}(B^{0}_{s}\to D^{-}_{s0}\ell^{\prime+}\nu_{\ell^{\prime}})$cdots $4.43\pm 1.54^{2q}$$8.84\pm 3.08^{4q}$ $\mathcal{B}(B^{0}_{s}\to D^{-}_{s2}\ell^{\prime+}\nu_{\ell^{\prime}})$cdots $2.72\pm 0.27$
$\mathcal{B}(B^{+}\to\overline{D}^{0}_{0}\tau^{+}\nu_{\tau})$cdots $5.23\pm 1.85$ $\mathcal{B}(B^{+}\to\overline{D}^{0}_{2}\tau^{+}\nu_{\tau})$cdots $2.15\pm 0.47$
$\mathcal{B}(B^{0}\to D^{-}_{0}\tau^{+}\nu_{\tau})$cdots $4.86\pm 1.72$ $\mathcal{B}(B^{0}\to D^{-}_{2}\tau^{+}\nu_{\tau})$cdots $1.97\pm 0.43$
$\mathcal{B}(B^{0}_{s}\to D^{-}_{s0}\tau^{+}\nu_{\tau})$cdots $6.75\pm 2.40^{2q}$$13.50\pm 4.80^{4q}$ $\mathcal{B}(B^{0}_{s}\to D^{-}_{s2}\tau^{+}\nu_{\tau})$cdots $1.73\pm 0.38$
@@
For the $B\to D_{0}\ell^{+}\nu_{\ell}$ decays, only two decay modes have been measured, and they are listed in the second column of Tab. 5, which are used to constrain the parameter $f^{BD_{0}}_{1}(0)$. Our constrained $f^{BD_{0}}_{1}(0)=0.38\pm 0.09$, which is obviously larger than $0.27\pm 0.03$ given in Ref. Verma:2011yw . Our branching ratio predictions of the semileptonic $B\to D_{0}\ell^{+}\nu_{\ell}$ decays are listed in the third column of Tab. 5. The branching ratio predictions of $B^{0}_{s}\to D^{-}_{s0}\ell^{+}\nu_{\ell}$ are different between the two quark state and the four quark state, the prediction with four quark state are 2 times of one with the two quark state. In the later analysis of $B\to DP\ell\nu_{\ell}$ with $D_{0}$ resonances, the results of the two quark state will be used.
@@
For the $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ decays, only $\mathcal{B}(B^{+}\to\overline{D}^{0}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ and $\mathcal{B}(B^{0}\to D^{-}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ have been measured, and they are listed in the fifth column of Tab. 5. We obtain $A_{1}^{BD^{*}_{2}}(0)=0.46\pm 0.06$ from two measured branching ratios, which are consistent with $0.63^{+0.11}_{-0.12}$ within $2\sigma$ errors given in Ref. Chen:2021ywv . Then one can predict the branching ratios of the semileptonic $B\to D^{*}\ell^{+}\nu_{\ell}$ decays, which are listed in the last column of Tab. 5. Decays $B^{0}_{s}\to D^{-}_{s2}\ell^{+}\nu_{\ell}$ have been calculated by the QCD sum rule approach for different scale parameter $\mu=2/3/4$ GeV Azizi:2014nta , for examples, $\mathcal{B}(B^{0}_{s}\to D^{-}_{s2}\ell^{\prime+}\nu_{\ell^{\prime}})=(3.07\pm 1.40)\times 10^{-3}$ and $\mathcal{B}(B^{0}_{s}\to D^{-}_{s2}\tau^{+}\nu_{\tau})=(1.03\pm 0.61)\times 10^{-3}$ for $\mu=4$ GeV. The predictions of $\mathcal{B}(B^{0}_{s}\to D^{-}_{s2}\ell^{\prime+}\nu_{\ell^{\prime}})$ in Ref. Azizi:2014nta are consistent with ours, nevertheless, the predictions of $\mathcal{B}(B^{0}_{s}\to D^{-}_{s2}\tau^{+}\nu_{\tau})$ in Ref. Azizi:2014nta are smaller than ours.
@@
Until now, the SU(3) flavor symmetry predictions of the three-body semileptonic $B\to D_{J}\ell^{+}\nu_{\ell}$ decays are quite coincident with their experimental data within $2\sigma$ errors. The SU(3) flavor breaking effects mainly come from different masses of $u$, $d$, and $s$ quarks. Since $m_{u,d}$ are much smaller than $m_{s}$, the SU(3) breaking effects due to a non-zero $m_{s}$ dominate the SU(3) breaking effects He:2014xha . If considering the SU(3) flavor breaking effects due to a non-zero $m_{s}$, the nonperturbative coefficients of the $B^{0}_{s}\to D^{-}_{s}/D^{*-}_{s}/D^{-}_{s0}/D^{*-}_{s2}\ell^{+}\nu_{\ell}$ decays are different from those of the $B^{+}\to\overline{D}^{0}/\overline{D}^{*0}/\overline{D}^{0}_{0}/\overline{D}^{*0}_{2}\ell^{+}\nu_{\ell}$ and $B^{0}\to D^{-}/D^{*-}/D^{-}_{0}/D^{*-}_{2}\ell^{+}\nu_{\ell}$ decays. As given in Tab. 5, decays $B^{0}_{s}\to D^{-}_{s}\mu^{+}\nu_{\mu}$ and $B^{0}_{s}\to D^{*-}_{s}\mu^{+}\nu_{\mu}$ have been measured, so one can estimate the SU(3) flavor breaking effects due to a non-zero $m_{s}$ in the $B\to D/D^{*}\ell^{+}\nu_{\ell}$ decays. Comparing our SU(3) flavor symmetry predictions and their experimental measurements of $B^{0}_{s}\to D^{-}_{s}/D^{*-}_{s}\mu^{+}\nu_{\mu}$ decays within $2\sigma$ errors, one can find the SU(3) breaking contributions to $\mathcal{B}(B^{0}_{s}\to D^{-}_{s}\ell^{+}\nu_{\ell})$ and $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s}\ell^{+}\nu_{\ell})$ should be less than roughly 23% and 20% of their experimental central values, respectively. After the $B^{0}_{s}\to D^{-}_{s0}/D^{*-}_{s2}\ell^{+}\nu_{\ell}$ decays are measured, one can estimate the SU(3) flavor breaking effects in the $B\to D_{0}/D^{*}_{2}\ell^{+}\nu_{\ell}$ decays.
III.2 Non-leptonic two-body $D_{J}\to DP$ decays

### v2

@@ III Decays $B\to DP\ell^{+}\nu_{\ell}$ with the $D^{*},D_{0},D^{*}_{2}$ resonances
For the $B\to DP\ell^{+}\nu_{\ell}$ decays with the resonances, when the decay widths of the resonance states are very narrow, the resonant branching ratios respect a simple factorization relation
@@
due to parity conservation, only $D_{J}=D_{0}/D^{*}/D^{*}_{2}$ resonances are considered for the $B\to DP\ell^{+}\nu_{\ell}$ decays Kim:1998np . And this result is also a good approximation for wider resonances. Eq. (27) will be used in our analysis for resonant $B\rightarrow D_{J}(\rightarrow DP)\ell^{+}\nu_{\ell}$ decays. We first calculate $\mathcal{B}(B\rightarrow D_{J}\ell^{+}\nu_{\ell})$ and $\mathcal{B}(D_{J}\rightarrow DP)$ by the SU(3) flavor symmetry, then obtain $\mathcal{B}(B\rightarrow D_{J}\ell^{+}\nu_{\ell},D_{J}\rightarrow DP)$. The width effects of some resonance states will be analyzed later.
III.1 Decays $B\to D_{J}\ell^{+}\nu_{\ell}$
The effective Hamiltonian is given in Eq. (8), and the amplitudes of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays also can be factorized in the product of the matrix elements of leptonic and hadronic currents. The differential branching ratios of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays are Ivanov:2019nqd
@@ III.1 Decays $B\to D_{J}\ell^{+}\nu_{\ell}$
${\cal H}_{U}$ $=$ $|H^{D_{J}}_{+}|^{2}+|H^{D_{J}}_{-}|^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\cal H}_{L}=|H^{D_{J}}_{0}|^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\cal H}_{S}=|H^{D_{J}}_{t}|^{2},$ (30)
where $\lambda\equiv\lambda(m_{B}^{2},m_{D_{J}}^{2},q^{2})$ with $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2ab-2ac-2bc$, and $m_{\ell}^{2}\leq q^{2}\leq(m_{B}-m_{D_{J}})^{2}$. The hadronic helicity amplitudes are defined through the hadronic matrix elements
$H^{D_{J}}_{m^{\prime}}$ $=$ $\epsilon^{*}_{\beta}(m^{\prime})\langle{D_{J}}(p,\varepsilon^{*})|\bar{c}\gamma^{\beta}(1-\gamma_{5})b|B(p_{B})\rangle,$ (31)
where $\varepsilon^{*}$ is the polarization of the $D^{*}$ and $D^{*}_{2}$ mesons, and $\epsilon_{\mu}(m)$ is the polarization of the virtual $W$ with $m=0,t,\pm 1$.
The hadronic helicity amplitudes related to the form factors are
@@
$H^{D^{*}_{2}}_{\pm}$ $=$ $\frac{2|\vec{p}_{D^{*}_{2}}|}{\sqrt{6}m_{D^{*}_{2}}}\left[(m_{B}+m_{D^{*}_{2}})A^{BD^{*}_{2}}_{1}(q^{2})\mp\frac{2m_{B}|\vec{p}_{D^{*}_{2}}|}{(m_{B}+m_{D^{*}_{2}})}V^{BD^{*}_{2}}(q^{2})\right],$ (37)
@@
for $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ decays, where $|\vec{p}_{D_{J}}|\equiv\sqrt{\lambda(m_{B}^{2},m_{D_{J}}^{2},q^{2})}/(2m_{B})$.
The form factors of the $B\to D_{J}$ transitions are defined as Cheng:2017pcq ; Cheng:2003sm
$D_{0}(p)\left|\bar{c}\gamma_{\mu}\gamma_{5}b\right|B(p_{B})\right>$ $=$ $-i\Bigg{[}f_{1}^{BD_{0}}(q^{2})\Big{(}(p+p_{B})_{\mu}-\frac{m_{B}^{2}-m_{D_{0}}^{2}}{q^{2}}q_{\mu}\Big{)}+f_{0}^{BD_{0}}(q^{2})\frac{m_{B}^{2}-m_{D_{0}}^{2}}{q^{2}}q_{\mu}\Bigg{]},$ (41)
$+iq_{\mu}(\varepsilon^{*}.p_{B})\frac{2m_{D^{*}}}{q^{2}}\Big{[}A^{BD^{*}}_{3}(q^{2})-A^{BD^{*}}_{0}(q^{2})\Big{]},$
$D^{*}_{2}(p,\varepsilon^{*})\left|\bar{c}\gamma_{\mu}(1-\gamma_{5})b\right|B(p_{B})\right>$ $=$ $\frac{2iV^{BD^{*}_{2}}(q^{2})}{m_{B}+m_{D^{*}_{2}}}\epsilon_{\mu\nu\alpha\beta}e^{*\nu}p^{\alpha}_{B}p^{\beta}$ (43)
$+2m_{D^{*}_{2}}\frac{e^{*}\cdot q}{q^{2}}q_{\mu}A^{BD^{*}_{2}}_{0}(q^{2})+(m_{B}+m_{D^{*}_{2}})\Big{(}e^{*}_{\mu}-\frac{e^{*}\cdot q}{q^{2}}q_{\mu}\Big{)}A^{BD^{*}_{2}}_{1}(q^{2})$
$-\frac{e^{*}\cdot q}{m_{B}+m_{D^{*}_{2}}}\Big{(}(p_{B}+p)_{\mu}-\frac{m_{B}^{2}-m^{2}_{D^{*}_{2}}}{q^{2}}q_{\mu}\Big{)}A^{BD^{*}_{2}}_{2}(q^{2}),$
where $s=q^{2}$ ($q=p_{B}-p$) and $e^{*\nu}\equiv\frac{\varepsilon^{*\mu\nu}\cdot p_{B\mu}}{m_{B}}$.
Now one can obtain the branching ratios of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays by the relevant form factors, which depend on the different methods. In this work, we use the SU(3) flavor symmetry to obtain the relations of the hadronic amplitudes, and the same relations are also true for the form factors. In terms of the SU(3) flavor symmetry, the hadronic helicity amplitudes defined in Eq. (31) can be parameterized as
$H^{D_{J}}_{m^{\prime}}=C^{D_{J}}_{0}B_{i}(D_{J})^{i},$ (44)
where $C^{D_{J}}_{0}$ are the non-perturbative coefficients under the SU(3) flavor symmetry. For the charmed four-quark states $D_{0}^{4q}$, $H^{D_{0}}_{m^{\prime}}=C^{4q,D_{0}}_{0}B_{i}(D_{J})^{ij}_{j}$. And the hadronic amplitude relations for the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays are summarized in Tab. 3.
Table 3: The hadronic amplitudes for $B\to D_{J}\ell^{+}\nu$ decays under the SU(3) flavor symmetry.
@@
Table 4: The experimental data and the SU(3) flavor symmetry predictions of $\mathcal{B}(B\to D/D^{*}\ell^{+}\nu_{\ell})$ within $2\sigma$ errors (in units of $10^{-2}$). adenotes that the experimental data are not used to constrain the parameter $C^{D^{*}}_{0}$.
@@
Table 5: The experimental data and the SU(3) flavor symmetry predictions of $\mathcal{B}(B\to D_{0}/D_{2}^{*}\ell^{+}\nu_{\ell})$ within $2\sigma$ errors. 2q(4q)denote the two(four) quark state predictions. $\mathcal{B}(B\to D_{0}/D^{*}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ is in unit of $10^{-3}$, and $\mathcal{B}(B\to D_{0}/D^{*}_{2}\tau^{+}\nu_{\tau})$ is in unit of $10^{-4}$.
Branching ratios Exp. data Belle-II:2022evt Our predictions Branching ratios Exp. data Belle-II:2022evt Our predictions
$\mathcal{B}(B^{+}\to\overline{D}^{0}_{0}\ell^{\prime+}\nu_{\ell^{\prime}})$ $4.2\pm 1.6$ $3.98\pm 1.30$ $\mathcal{B}(B^{+}\to\overline{D}^{*0}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ $2.9\pm 0.6$ $3.20\pm 0.30$
$\mathcal{B}(B^{0}\to D^{-}_{0}\ell^{\prime+}\nu_{\ell^{\prime}})$ $3.9\pm 1.4$ $3.71\pm 1.21$ $\mathcal{B}(B^{0}\to D^{*-}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ $2.7\pm 0.6$ $2.99\pm 0.29$
$\mathcal{B}(B^{0}_{s}\to D^{-}_{s0}\ell^{\prime+}\nu_{\ell^{\prime}})$cdots $4.43\pm 1.54^{2q}$$8.84\pm 3.08^{4q}$ $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\ell^{\prime+}\nu_{\ell^{\prime}})$cdots $2.72\pm 0.27$
$\mathcal{B}(B^{+}\to\overline{D}^{0}_{0}\tau^{+}\nu_{\tau})$cdots $5.23\pm 1.85$ $\mathcal{B}(B^{+}\to\overline{D}^{*0}_{2}\tau^{+}\nu_{\tau})$cdots $2.15\pm 0.47$
$\mathcal{B}(B^{0}\to D^{-}_{0}\tau^{+}\nu_{\tau})$cdots $4.86\pm 1.72$ $\mathcal{B}(B^{0}\to D^{*-}_{2}\tau^{+}\nu_{\tau})$cdots $1.97\pm 0.43$
$\mathcal{B}(B^{0}_{s}\to D^{-}_{s0}\tau^{+}\nu_{\tau})$cdots $6.75\pm 2.40^{2q}$$13.50\pm 4.80^{4q}$ $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\tau^{+}\nu_{\tau})$cdots $1.73\pm 0.38$
@@
For the $B\to D_{0}\ell^{+}\nu_{\ell}$ decays, only two decay modes have been measured, and they are listed in the second column of Tab. 5, which are used to constrain the parameter $f^{BD_{0}}_{1}(0)$. Our constrained $f^{BD_{0}}_{1}(0)=0.38\pm 0.09$, which is obviously larger than $0.27\pm 0.03$ given in Ref. Verma:2011yw . Our branching ratio predictions of the $B\to D_{0}\ell^{+}\nu_{\ell}$ decays are listed in the third column of Tab. 5. The branching ratio predictions of $B^{0}_{s}\to D^{-}_{s0}\ell^{+}\nu_{\ell}$ are different between the two quark state and the four quark state, the prediction with four quark state are 2 times of one with the two quark state. In the later analysis of $B\to DP\ell\nu_{\ell}$ with $D_{0}$ resonances, the results of the two quark state will be used.
@@
For the $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ decays, only $\mathcal{B}(B^{+}\to\overline{D}^{*0}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ and $\mathcal{B}(B^{0}\to D^{*-}_{2}\ell^{\prime+}\nu_{\ell^{\prime}})$ have been measured, and they are listed in the fifth column of Tab. 5. We obtain $A_{1}^{BD^{*}_{2}}(0)=0.46\pm 0.06$ from two measured branching ratios, which are consistent with $0.63^{+0.11}_{-0.12}$ within $2\sigma$ errors given in Ref. Chen:2021ywv . The branching ratio predictions of the $B\to D^{*}_{2}\ell^{+}\nu_{\ell}$ decays are listed in the last column of Tab. 5. Decays $B^{0}_{s}\to D^{*-}_{s2}\ell^{+}\nu_{\ell}$ have been calculated by the QCD sum rule approach for different scale parameter $\mu=2/3/4$ GeV Azizi:2014nta , for examples, $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\ell^{\prime+}\nu_{\ell^{\prime}})=(3.07\pm 1.40)\times 10^{-3}$ and $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\tau^{+}\nu_{\tau})=(1.03\pm 0.61)\times 10^{-3}$ for $\mu=4$ GeV. The predictions of $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\ell^{\prime+}\nu_{\ell^{\prime}})$ in Ref. Azizi:2014nta are consistent with ours, nevertheless, the predictions of $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s2}\tau^{+}\nu_{\tau})$ in Ref. Azizi:2014nta are smaller than ours.
@@
Until now, most of the SU(3) flavor symmetry predictions of the $B\to D_{J}\ell^{+}\nu_{\ell}$ decays are quite coincident with their experimental data within $2\sigma$ errors. The SU(3) flavor breaking effects mainly come from different masses of $u$, $d$, and $s$ quarks. Since $m_{u,d}$ are much smaller than $m_{s}$, the SU(3) breaking effects due to a non-zero $m_{s}$ dominate the SU(3) breaking effects He:2014xha . If considering the SU(3) flavor breaking effects due to a non-zero $m_{s}$, the non-perturbative coefficients of the $B^{0}_{s}\to D^{-}_{s}/D^{*-}_{s}/D^{-}_{s0}/D^{*-}_{s2}\ell^{+}\nu_{\ell}$ decays are different from those of the $B^{+}\to\overline{D}^{0}/\overline{D}^{*0}/\overline{D}^{0}_{0}/\overline{D}^{*0}_{2}\ell^{+}\nu_{\ell}$ and $B^{0}\to D^{-}/D^{*-}/D^{-}_{0}/D^{*-}_{2}\ell^{+}\nu_{\ell}$ decays. As given in Tab. 5, decays $B^{0}_{s}\to D^{-}_{s}\mu^{+}\nu_{\mu}$ and $B^{0}_{s}\to D^{*-}_{s}\mu^{+}\nu_{\mu}$ have been measured, so one can estimate the SU(3) flavor breaking effects due to a non-zero $m_{s}$ in the $B\to D/D^{*}\ell^{+}\nu_{\ell}$ decays. Comparing our SU(3) flavor symmetry predictions and their experimental measurements of $B^{0}_{s}\to D^{-}_{s}/D^{*-}_{s}\mu^{+}\nu_{\mu}$ decays within $2\sigma$ errors, one can find the SU(3) breaking contributions to $\mathcal{B}(B^{0}_{s}\to D^{-}_{s}\ell^{+}\nu_{\ell})$ and $\mathcal{B}(B^{0}_{s}\to D^{*-}_{s}\ell^{+}\nu_{\ell})$ should be less than roughly 23% and 20% of their experimental central values, respectively. After the $B^{0}_{s}\to D^{-}_{s0}/D^{*-}_{s2}\ell^{+}\nu_{\ell}$ decays are measured, one can estimate the SU(3) flavor breaking effects in the $B\to D_{0}/D^{*}_{2}\ell^{+}\nu_{\ell}$ decays.
III.2 Decays $D_{J}\to DP$